I need help with three math problems:

#1.) 4x^2 + 10x +6

#2.) x^2 - 5x+6 = 0

#3.) (6x + 2) (3x - 12) = 0

Thanks!

Answer by mananth(12270)

#1.) 4x^2 + 10x +6
a=4, b=10 c= 6
ac= 24
4x^2+6x+4x+6
2x(2x+3)+2(2x+3)
(2x+2)(2x+3)
2(x+2)(2x+3)



#2.) x^2 - 5x+6 = 0
x^2-3x-2x+6
x(x-3)-2(x-3)
(x-3)(x-2)=0
x=3 OR 2

#3.) (6x + 2) (3x - 12) = 0
If if you have to find the roots
(6x+2)=0
6x=-2
x= -2/6
x=-1/3
OR
3x-12=0
3x=12
x=4